3 Solve the initial value problem using Laplace Transforms: y - y = 0, y(0) = 1, y (O) = 0, y (0) = 1, y(0) = 0 b) y+9y =
Answer #1

dithemtad equotrun C3) (a) Givm = 0, (0)1, acoJ L- L[O] L[ -La ye )(o)-aco) -LJ=o sY LE 0 -s -0 -LLCa] s Nw A s2 S- (SXS)) ts +e ) (2+1)(s-1 ) +D(s+1) (s1) A (S-)( Att2(-Atc+3) +s (A-8 A c+D Cmporsing the t (2) A -B+De0 (3)Similandy, Bs +C S-I s+l (S+)) (S-1 ) (s+1) s (A+B+D) + s? ( -A+s+D) +s ( A-R+D) -A-C +D Comporong hoth side, ue get (1) -(2Sen +2 Cos 2 Cos e , ao TauknLoplac e h both side, we ged LCJ -L[e L (S-1)9) Now Bs (Ppanka frachm) + (s49) S-I AC5249)Bst*Addery 2) $(3), we get 14) 9A - 13= | we get Addeng 10 A - - to .B 10 10 So V10 (s-1) (s?+9) s?+ 9 Hence to + to L[a ] s?+9